0.890 atm and a telnperature o135.0C. 8. Calculate the volume of 25.0 g of the following liquids: acetone (d = 0.792 g/mL) chloroform (d 1.49 g/mL) a. b.
Answer #1

density = mass / volume

0.792 g/mL = 25.0 g / volume

volume = 25.0 g / 0.792 g/mL = 31.6 mL

Therefore, the volume = 31.6 mL

b) density = mass / volume

1.49 g/mL = 25.0 g / volume

volume = 25.0 g / 1.49 g/mL = 16.8 mL

Therefore, the volume = 16.8 mL